# Enhancing Adversarial Robustness of Vision Language Models via Adversarial Mixture Prompt Tuning

### v1

@@ 5.4 Ablation Study
To verify the effectiveness of AMPT, we conduct a set of ablation studies. We conduct the experiment in the Caltech101 dataset with the 16-shot training setting with class-specific context prompts.
Effects of Different Components. Here we try to conduct ablation experiments on various components of the experiment. We use the single adversarial cue fine-tuning method as the baseline and add adversarial mixture prompts without adding the conditional prompt weight router. Furthermore, we add our conditional prompt weight router. The experimental results are shown in Table 5.
@@
Component Clean PGD
Baseline 86.29 56.75
Baseline+Mixture 87.14 59.19
Baseline+Mixture+Router 87.01 60.12
The results can verify the effectiveness of each component in our AMPT. Only adding mixture prompts but without weight router can bring a 2.44% improvement compared with the baseline against PGD attack, while conditional prompt weight router can further bring a 0.93% improvement. This result shows that the feature diversity brought by adversarial mixture prompts and the adaptability brought by the conditional prompt weight router work together to improve the generalization of the VLMs to deal with adversarial examples.
Table 6: Ablation Study towards Selection of Prompt Number.
Prompt number Clean PGD
1 86.29 56.75
3 86.98 59.55
5 87.66 59.67
8 87.01 60.12
10 86.98 60.04
Selection of Prompt number. We explore the selection of prompt numbers. We select the following text prompt number of our AMPT as 1, 3, 5, 8, 10, 12, and the result can be viewed in Table 6. From the results, when the prompt number increases at the beginning (from number 1 to 3), the adversarial robustness of AMPT will obviously increase. However, when it further increases (from number 8 to 10), the robustness remains basically unchanged. It can be explained that as the prompt number increases, the difficulty of prompt optimization also increases. Thus, we select the prompt number 8 as the default setting. Considering the calculation overhead, we also select prompt number 3 for the partial dataset.
Table 7: Ablation Study towards Selection of Hyper-parameter $\tau$.
$\tau$ Clean PGD
0.3 86.94 59.59
0.5 87.26 59.71
0.7 87.01 60.12
0.8 86.97 60.04
1 87.14 59.67
Selection of Hyper-parameter $\tau$. The temperatures $\tau$ can control the adjustment strength of the conditional prompt weight router. While smaller $\tau$ means larger adjustment strength, larger $\tau$ means smaller adjustment strength. We select the following text prompt numbers of our AMPT as 0.3, 0.5, 0.7, 0.8, and 1, and the results can be found in Table 7. Based on the experimental results, we select the Hyper-parameter $\tau$ to 0.7 in our final setting.
6 Limitation
From the perspective of computational cost, as a parameter-efficient method, AMPT is lightweight and can be quickly optimized. However, our AMPT tuning introduces additional parameters and still has more computational cost during the training process. In addition, similar to the other parameter-efficient method, AMPT relies on pre-trained VLMs with pretty initial adversarial robustness, which needs to be further explored in the future.
@@ 7 Conclusion
In this work, we focused on the overfitting problem of adversarial prompt tuning, and found that simply increasing the length of the text prompt led to the learning difficulty while increasing the prompt number was more likely to improve the adversarial robustness of the VLMs. Based on the observation, we proposed an adversarial prompt tuning method named Adversarial Mixture Prompt Tuning (AMPT), which introduced adversarial mixture prompts to obtain more general text features to improve adversarial robustness. In addition, we proposed a conditional prompt weight router, which generated different prompt weights for different text prompts based on the characteristics of the input image to improve the adaptability of adversarial mixture prompts. Meanwhile, we theoretically proved the effectiveness of the conditional prompt weight router. A series of experimental results showed that our AMPT could achieve obvious improvements for the in-distribution robustness on different datasets and had well-transfer robustness across different datasets. In the future, AMPT can transform into different modals, e.g., visual prompts, and will be applied to more different downstream tasks.
Appendix A Appendix
A.1 The proof of Theorem 1
Here we assume each text prompt keeps unchanged and only optimize the prompt weights $\tilde{w}=\{\tilde{w}_{1},\tilde{w}_{2},\cdots,\tilde{w}_{K}\}$ from initial weights $\tilde{w}^{0}=\{1/K,1/K,\cdots,1/K\}$. Then as for any adversarial example $\tilde{x}^{{}^{\prime}}$, without loss generalization, we can assmue a corresponding text prompt $t^{k}$ exist error risk as follows:
$0<\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{k},y)<\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{1},y)<\cdots<\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{k-1},y)<\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{k+1},y)<\cdots<\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{K},y).$ (13)
Thus based on the Gradient Descent operation, the expectation of corresponding weights can be formulated as follows:
$\mathbb{E}(\tilde{w}_{k}^{{}^{\prime}})=\frac{1}{K}-\eta\cdot\frac{\partial\frac{1}{K}\sum_{k}^{K}\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{k},y)}{\partial\tilde{w}_{k}},$ (14)
and here we assume that different weights are independent of each other, so we can obtain the $\mathbb{E}(\tilde{w}_{K}^{{}^{\prime}})$ as follows:
$\mathbb{E}(\tilde{w}_{k}^{{}^{\prime}})=\frac{1}{K}-\eta\cdot\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{k},y).$ (15)
Then, we apply softmax to each $\mathbb{E}(\tilde{w}_{k}^{{}^{\prime}})$, and denote the result of applying softmax to $\mathbb{E}(\tilde{w}_{k}^{{}^{\prime}})$ is $\mathbb{E}(\tilde{w}_{k})$, then we get:
$\mathbb{E}(\tilde{w}_{k}^{{}^{\prime}})=\frac{e^{\frac{1}{K}-\eta\cdot\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{k},y)}}{\sum_{j}^{K}e^{\frac{1}{K}-\eta\cdot\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{j},y)}}.$ (16)
For simplicity, we let $\mathcal{R}_{k}=\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{k},y)$, so Eq. (16)It can be further simplified to:
$\mathbb{E}(\tilde{w}_{k})=\frac{e^{-\eta\cdot\mathcal{R}_{k}}}{\sum_{j}^{K}e^{-\eta\cdot\mathcal{R}_{j}}}.$ (17)
Then the entire error expectation of adversarial examples can be formulated as follows:
$\mathbb{E}(\sum_{k}^{K}\tilde{w}_{k}\mathcal{R}_{k})=\mathbb{E}(\frac{1}{K}\sum_{k}^{K}\mathcal{R}_{k})+\sum_{k}^{K}(\mathbb{E}(\tilde{w}_{k})-\frac{1}{K})*\mathcal{R}_{k}.$ (18)
After substituting $\mathbb{E}(\tilde{w}_{k})$ into $\sum_{k}^{K}(\mathbb{E}(\tilde{w}_{k})-\frac{1}{K})*\mathcal{R}_{k}$, we have:
$\sum_{k}^{K}(\mathbb{E}(\tilde{w}_{k})-\frac{1}{K})*\mathcal{R}_{k}=\sum_{k}^{K}(\frac{e^{-\eta\cdot\mathcal{R}_{k}}}{\sum_{j}^{K}e^{-\eta\cdot\mathcal{R}_{j}}}-\frac{1}{K})\cdot\mathcal{R}_{k}.$ (19)
We assume the following relation holds:
$\sum_{k}^{K}(\frac{e^{-\eta\cdot\mathcal{R}_{k}}}{\sum_{j}^{K}e^{-\eta\cdot\mathcal{R}_{j}}}-\frac{1}{K})\cdot\mathcal{R}_{k}<0,$ (20)
the above inequality is equivalent to the following inequality:
$\sum_{k}^{K}(e^{-\eta\cdot\mathcal{R}_{k}}-\frac{\sum_{j}^{K}e^{-\eta\cdot\mathcal{R}_{j}}}{K})\cdot\mathcal{R}_{k}<0.$ (21)
When $K=2$, without loss of generality, we assume that $\mathcal{R}_{1}<\mathcal{R}_{2}$, then we have:
$\sum_{k}^{K}(e^{-\eta\cdot\mathcal{R}_{k}}-\frac{\sum_{j}^{K}e^{-\eta\cdot\mathcal{R}_{j}}}{K})\cdot\mathcal{R}_{k}=\frac{e^{-\eta\cdot\mathcal{R}_{1}}-e^{-\eta\cdot\mathcal{R}_{2}}}{2}\cdot(\mathcal{R}_{1}-\mathcal{R}_{2}).$ (22)
Since $\mathcal{R}_{1}<\mathcal{R}_{2}$, So we have $\mathcal{R}_{1}-\mathcal{R}_{2}<0$ and $\frac{e^{-\eta\cdot\mathcal{R}_{1}}-e^{-\eta\cdot\mathcal{R}_{2}}}{2}>0$, which means when $K=2$, the assumption holds. Next, we assume that the assumption holds when $K=n$, and when $K=n+1$, without loss of generality, suppose $\mathcal{R}_{n+1}$ satisfies:
$\mathcal{R}_{n+1}<\mathcal{R}_{1}<\dots<\mathcal{R}_{n},$ (23)
then we have:
$\sum_{k}^{n+1}(e^{-\eta\cdot\mathcal{R}_{k}}-\frac{\sum_{j}^{n+1}e^{-\eta\cdot\mathcal{R}_{j}}}{K})\cdot\mathcal{R}_{k}$ $=\sum_{k}^{n+1}e^{-\eta\cdot\mathcal{R}_{k}}\cdot\mathcal{R}_{k}-\frac{\sum_{j}^{n+1}e^{-\eta\cdot\mathcal{R}_{j}}}{n+1}\cdot\sum_{k}^{n+1}\cdot\mathcal{R}_{k}.$ (24)
By the induction hypothesis, we have:
$\sum_{k}^{n+1}e^{-\eta\cdot\mathcal{R}_{k}}\cdot\mathcal{R}_{k}$ $=\sum_{k}^{n}e^{-\eta\cdot\mathcal{R}_{k}}\cdot\mathcal{R}_{k}+e^{-\eta\cdot\mathcal{R}_{n+1}}\cdot\mathcal{R}_{n+1}$ (25)
$<\frac{\sum_{j}^{n}e^{-\eta\cdot\mathcal{R}_{j}}}{n}\cdot\sum_{k}^{n}\mathcal{R}_{k}+e^{-\eta\cdot\mathcal{R}_{n+1}}\cdot\mathcal{R}_{n+1},$
meanwhile, we can have:
$\frac{\sum_{j}^{n+1}e^{-\eta\cdot\mathcal{R}_{j}}}{n+1}\cdot\sum_{k}^{n+1}\cdot\mathcal{R}_{k}$ $=\frac{\sum_{j}^{n}e^{-\eta\cdot\mathcal{R}_{j}}}{n+1}\cdot\sum_{k}^{n}\mathcal{R}_{k}+\frac{\sum_{j}^{n}e^{-\eta\cdot\mathcal{R}_{j}}}{n+1}\cdot\mathcal{R}_{n+1}$ (26)
$+\frac{e^{-\eta\cdot\mathcal{R}_{n+1}}}{n+1}\cdot\sum_{k}^{n}\mathcal{R}_{k}+\frac{e^{-\eta\cdot\mathcal{R}_{n+1}}}{n+1}\cdot\mathcal{R}_{n+1}.$
Substituting Eq. (26) and Eq. (25) into Eq. (24), we can obtain:
$\sum_{k}^{n+1}(e^{-\eta\cdot\mathcal{R}_{k}}-\frac{\sum_{j}^{n+1}e^{-\eta\cdot\mathcal{R}_{j}}}{n+1})\cdot\mathcal{R}_{k}$ $<[(\frac{\sum_{j}^{n}e^{-\eta\cdot\mathcal{R}_{j}}}{n(n+1)}\cdot\sum_{k}^{n}\mathcal{R}_{k}-\frac{e^{-\eta\cdot\mathcal{R}_{n+1}}}{n+1}\cdot\sum_{k}^{n}\mathcal{R}_{k})$ (27)
$+(\frac{n\cdot e^{-\eta\cdot\mathcal{R}_{n+1}}}{n+1}\cdot\mathcal{R}_{n+1}-\frac{\sum_{j}^{n}e^{-\eta\cdot\mathcal{R}_{j}}}{n+1}\cdot\mathcal{R}_{n+1})]$
$=\frac{1}{n+1}\cdot(\frac{\sum_{j}^{n}e^{-\eta\cdot\mathcal{R}_{j}}}{n}-e^{-\eta\cdot\mathcal{R}_{n+1}})\cdot(\sum_{k}^{n}\mathcal{R}_{k}-n\cdot\mathcal{R}_{n+1}).$
Since we have previously assumed that $\mathcal{R}_{n+1}<\mathcal{R}_{1}<\dots<\mathcal{R}_{n}$, so we can deduce that:
$\sum_{k}^{n}\mathcal{R}_{k}-n\cdot\mathcal{R}_{n+1}>0,$ (28)
and:
$\frac{\sum_{j}^{n}e^{-\eta\cdot\mathcal{R}_{j}}}{n}-e^{-\eta\cdot\mathcal{R}_{n+1}}<0.$ (29)
Therefore, we can conclude that:
$\sum_{k}^{n+1}(e^{-\eta\cdot\mathcal{R}_{k}}-\frac{\sum_{j}^{n+1}e^{-\eta\cdot\mathcal{R}_{j}}}{K})\cdot\mathcal{R}_{k}<0.$ (30)
Therefore, by mathematical induction, we can conclude that the following holds for any K:
$\sum_{k}^{K}(e^{-\eta\cdot\mathcal{R}_{k}}-\frac{\sum_{j}^{K}e^{-\eta\cdot\mathcal{R}_{j}}}{K})\cdot\mathcal{R}_{k}<0.$ (31)
Substituting Eq. (31) into Eq. (18), we obtain:
$\mathbb{E}(\sum_{k}^{K}\tilde{w}_{k}\mathcal{R}_{k})<\mathbb{E}(\frac{1}{K}\sum_{k}^{K}\mathcal{R}_{k}).$ (32)
Then the Theorem 1 is proved.
A.2 Details on Generalization cross Different Datasets
As shown in Table 8, we present the robust generalization performance across different target datasets when trained on Caltech101 as the source domain. The results show that our AMPT-UC consistently outperforms APT-UC in adversarial robustness across 8 datasets, and achieves higher accuracy on 6 datasets, indicating superior generalization ability and enhanced resilience against diverse attacks.
Table 8: Detailed results comparing the Robust Generalization (%) between APT-UC and AMPT-UC cross 9 different datasets based on Caltech101 adversarial prompts.
Method Metric OxfordPets Flowers102 Cars FGVC DTD SUN397 Food101 EuroSAT UCF101 Average Clean 29.95 14.01 8.10 1.83 16.84 14.19 15.52 12.81 19.51 14.75 PGD 10.06 3.17 0.87 0.36 7.74 2.25 1.43 1.14 3.49 3.39 APT-UC AA 8.67 2.15 0.39 0.33 6.74 1.63 0.77 0.58 2.72 2.66 Clean 45.08 16.00 13.01 2.34 16.02 20.17 15.3 11.43 26.20 18.39 PGD 9.10 4.02 1.02 0.81 7.68 4.10 2.24 10.67 4.34 4.89 MAPT-UC AA 7.14 2.84 0.53 0.75 6.97 4.02 1.31 10.57 3.09 4.14
A.3 Trade-off between Accuracy and Robustness
As shown in Fig. 3, we compare the performance improvement per dataset of our adversarially-trained prompt over the standard-trained prompt for unified context. Most adversarially trained vision models tend to improve robustness at the expense of accuracy, and adversarially trained prompts also exhibit this trade-off, which is expected. More importantly, we observe that for most datasets, the gain in robustness outweighs the drop in accuracy. Specifically, AMPT improves adversarial robustness by an average of +7.60%, while incurring only a modest drop of -3.03% in accuracy. For instance, on OxfordPets, robustness increases significantly by +8.61%, with a slight gain of +0.23% in accuracy. These results suggest that our method achieves a relatively favorable trade-off between accuracy and robustness.
Figure 3: Trade-off between Accuracy and Robustness ($M=16$).

### v2

@@ 5.4 Ablation Study
To verify the effectiveness of MoAPT, we conduct a set of ablation studies. We conduct the experiment in the Caltech101 with the 16-shot setting. All the setting keep the same with the default setting if without additional instructions.
@@
Component Clean PGD AA
Baseline 86.29 56.75 53.43
Baseline+Mixture 87.06 57.36 54.60
Baseline+Mixture+Router 87.14 57.69 55.05
Effects of Different Components. We conduct ablation studies on different components. Starting from the single adversarial cue fine-tuning baseline, we first add adversarial mixture prompts, and then further incorporate the conditional prompt weight router. Table 4 reports the results on Caltech101, with others given in Appendix 1.2.
The results confirm the contribution of each module in MoAPT. Introducing mixture prompts without the weight router yields a modest robustness gain over the baseline, while integrating the conditional prompt-weight router provides a further improvement of 1.06%/1.19% under PGD/AA attacks and a 0.99% increase in clean accuracy. These findings indicate that the feature diversity introduced by adversarial mixture prompts and the adaptive weighting enabled by the conditional router work in a complementary manner, jointly enhancing both the robustness and generalization of VLMs.
Table 5: Ablation Study towards Prompt Number.
Prompt number Clean PGD AA
1 86.29 56.75 53.43
2 86.13 56.98 53.66
4 86.69 57.45 54.52
6 87.22 57.04 54.32
8 87.14 57.69 55.05
10 87.34 56.80 54.40
12 87.55 57.20 54.44
Selection of Prompt number. We explore the selection of prompt numbers. We select the following text prompt number of our MoAPT as 1, 2, 4, 6, 8, 10, 12, and the result can be viewed in Table 5. From the results, when the number of prompts increases at the beginning (from number 1 to 8), the adversarial robustness of MoAPT will obviously increase. However, when it further increases (from number 8 to 12), the robustness remains basically unchanged. It can be explained that as the number of prompts increases, the difficulty of prompt optimization also increases. Thus, we select the prompt number 8 as the default setting.
Table 6: Ablation Study towards Hyper-parameter $\tau$.
$\tau$ Clean PGD AA
0.3 85.72 55.98 53.18
0.5 86.86 57.93 54.32
0.7 87.14 57.69 55.05
0.9 87.14 57.32 54.69
1.1 87.05 57.77 54.56
Selection of Hyper-parameter $\tau$. The temperatures $\tau$ can control the adjustment strength of the conditional prompt weight router. While smaller $\tau$ means larger adjustment strength, larger $\tau$ means smaller adjustment strength. We select the following $\tau$ of MoAPT as 0.3, 0.5, 0.7, 0.9, and 1.1, and the results can be found in Table 6. Based on the experimental results, we select the Hyper-parameter $\tau$ to 0.7.
5.5 Computational Cost
Despite adding multiple prompts, MoAPT still remains a parameter-efficient and highly competitive method as shown in Table 7. The inference memory and time costs of MoAPT are slightly higher than those of APT but are still lower than those of FAP, indicating that it maintains high inference efficiency while ensuring robustness.
Table 7: Calculation Overhead. The results are conducted based on RTX 4090 in 16-shot setting of each epoch with Caltech101.
Method VPT APT FAP MoAPT
Training Memory Cost 6730M 2798M 4204M 14384M
Training Time Cost 30s 14s 165s 60s
Testing Memory Cost 2246M 5626M 7478M 5838M
Testing Time Cost 5.00s 8.42s 14.01s 10.66s
@@ 6 Conclusion
In this work, we focused on the overfitting problem of adversarial prompt tuning, and found that simply increasing the length of the text prompt led to the learning difficulty while increasing the number of prompts was more likely to improve the adversarial robustness of the VLMs. Based on the observation, we propose Mixture of Adversarial Prompt Tuning (MoAPT), which introduces adversarial mixture prompts to obtain more general text features, and proposes a conditional prompt weight router to further improve the adaptability of adversarial mixture prompts. Our theoretical analysis validates the effectiveness of the router. Extensive experiments demonstrate that MoAPT consistently improves in-distribution robustness and exhibits strong transfer robustness across diverse datasets.